# Predicting topological invariants and unconventional superconducting pairing from density of states and machine learning

### v1

@@ VI Conclusions
Importantly, the ML models correctly predict $B=0$ for trivial edge modes. Therefore, ML can be an important tool to distinguish between topologically trivial and nontrivial edge modes when the local density of states at zero energy is obtained from scanning tunneling spectroscopic experiments. This method can also distinguish between topological phases with different numbers of chiral edge states. We stress that the ML model trained with clean superconductors has a fair prediction of the topological phase diagram of disordered systems. Similarly, ML algorithms trained with disordered theoretical models may be able to predict the topological invariant of a real superconducting sample, which is naturally more complex.
@@ VII Acknowledgements
We thank Eric Andrade and Diego Ferreira for helpful discussions. This work was supported by the National Council for Scientific and Technological Development – CNPq (F.N., A.C., R.C., R.G.P.) and by a grant from the Simons Foundation (Grant No. 1023171, R.C., R.G.P.). Research at IIP-UFRN is supported by Brazilian ministries MEC and MCTI. We also acknowledge the High-Performance Computing Center (NPAD) at UFRN for providing computational resources. AC acknowledges the partial financial support by the Alagoas State Research Agency (FAPEAL) (Grant No. APQ2022021000153202202100015320220210001532022021000153).

### v2

@@ VI Conclusions
Importantly, the ML models correctly predict $B=0$ for trivial edge modes. Therefore, ML can be an important tool to distinguish between topologically trivial and nontrivial edge modes when the local density of states at zero energy is obtained from scanning tunneling spectroscopic experiments. This method can also distinguish between topological phases with different numbers of chiral edge states. We stress that the ML model trained with clean superconductors has a fair prediction of the topological phase diagram of disordered systems. Similarly, ML algorithms trained with disordered theoretical models may be able to predict the topological invariant of a real superconducting sample, which is naturally more complex. However, before applying the method to real world data, it would be important to generalize our approach and train ML models with different geometries, other kinds of disorder, and model parameters that are close to the experimental setting.
Figure 9: Weights $W$ in the linear function $Z=WX+b$ associated with the LDOS for every lattice site. We display the weights for the phases with $B=-1$ (top left), $B=0$ (top right), and $B=2$ (bottom left). The sum of weights for $B=-1$ and $B=2$ is shown in the bottom right.
@@ VII Acknowledgements
We thank Eric Andrade and Diego Ferreira for helpful discussions. This work was supported by the National Council for Scientific and Technological Development – CNPq (F.N., A.C., R.C., R.G.P.) and by a grant from the Simons Foundation (Grant No. 1023171, R.C., R.G.P.). Research at IIP-UFRN is supported by Brazilian ministries MEC and MCTI. We also acknowledge the High-Performance Computing Center (NPAD) at UFRN for providing computational resources. AC acknowledges partial financial support by the Alagoas State Research Agency (FAPEAL) (Grant No. APQ2022021000153202202100015320220210001532022021000153).
Appendix A Weights in the ML model
To gain insight into how the ML model extracts features from the LDOS to classify the phases, here we discuss the weights in the neural network for the simple case of Fig. 4, where we used no hidden layer. The exit layer has three nodes, which correspond to each of the classes $B=-1,0,+2$. Each node computes the linear function $Z=WX+b$, where $X$ is an array containing the values of zero-energy LDOS for all lattice sites, $W$ is an array with the weights of the neural network, and $b$ is a bias scalar. Both $W$ and $b$ are estimated during the training of the ML model. The value of $Z$ is then transformed to a probability through the softmax activation function. In Fig. 4, the computed values of $b$ are $b=0.9625709,2.04757,-2.566742$ for $B=-1,0,+2$, respectively. The weights $W$, computed for each $B$, are shown in Fig. 9 for each position of the lattice. One can see that if the LDOS is higher in the bulk, the value of $Z$ for $B=0$ is also high, leading to a high probability that the sample has $B=0$. If, however, the LDOS is higher along the edges, the $Z$ values for $B=-1,+2$ increase. Interestingly, the weights for $B=-1,+2$ display some structure along the diagonals near the corners. The last plot shows the sum of $W$ for $B=-1$ with $W$ for $B=2$. In this case the diagonal structure disappears. This means that the structure that the neural network expects in the diagonals for $B=-1$ is complementary to the one for $B=+2$. We stress, however, that this analysis holds for a neural network without hidden layers. The presence of two hidden layers can change the expected structure of the LDOS for each class. Nevertheless, this analysis illustrates how the edge states and structures near the corners can be used to predict the Bott index.